OpenStudy (anonymous):

Work and power are different because work involves force and distance while power involves force and velocity. work involves distance and mass while power involves distance and time. work involves energy and time while power involves force and mass. work involves velocity and time while power involves energy and force.

OpenStudy (anonymous):

Work is force (force) times distance (d). Power (p) is work over time (t).If you substitute force times distance for work, it simplifies to force times velocity (v) because distance over time is velocity. like this: w=fd p=w/t=fd/t=f(d/t)=fv So work involves force and distance and power involves force and velocity